solve the system:
4x+5y = -24
x-5y = 19

Answer by checkley71(8403)

4X+5Y=-24
X-5Y=19 NOW ADD THESE TWO EQUATIONS
---------------------
5X=-5
X=-5/5
X=-1
4*-1+5Y=-24
-4+5Y=-24
5Y=-24+4
5Y=-20
Y=-20/5
Y=-4
PROOF
-1-5*-4=19
-1+20=19
19=19